40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 6 22:47:41 2021, gautam, Update, ALS, Noisy ALS ALSX_excess.pngbudget.pdf
    Reply  Thu Jan 7 19:16:11 2021, gautam, Update, ALS, Noisy ALS ALS_ERR_MON.pdfAUXnoise.pdf
       Reply  Thu Jan 7 20:07:27 2021, Koji, Update, ALS, Noisy ALS 
          Reply  Thu Jan 7 21:16:22 2021, gautam, Update, ALS, Noisy ALS 
             Reply  Thu Jan 7 23:25:19 2021, Koji, Update, ALS, Noisy ALS 
          Reply  Fri Jan 8 20:01:11 2021, gautam, Update, ALS, Noisy ALS ALSX_IR_green.pdf
Message ID: 15752     Entry time: Thu Jan 7 19:16:11 2021     In reply to: 15751     Reply to this: 15753
Author: gautam 
Type: Update 
Category: ALS 
Subject: Noisy ALS 

I'm also wondering why the error monitors for the X and Y loops report such wildly different spectra for the suppressed frequency noise of the AUX laser relative to the cavity length, see Attachment #1. The y-axis should be approximately Hz/rtHz. In both cases, the servo's error point monitor is connected to the DAQ system via a G=10 SR560. With the SR785, I measure for EX a nice bucket-shaped spectrum, bottoming out at ~10 uV/rtHz around 40 Hz, see Attachment #2. The SR560 should have an input-referred noise much less than this at the G=10 setting. The ADC noise level is only ~5 uV/rtHz, and indeed, the EY spectrum shows the expected shape. So what's up with the EX error mon? Tried swapping out the SR560 for a different unit, no change. And both the SR560 noise, and the ADC noise, check out when everything is checked individually. So some kind of interaction once everything is connected together, but it's only present at EX...

Today, I tried switching the EX and EY fibers going into the beat mouth, but I preserved the channel mapping after the beat mouth by switching the electrical outputs as well (the goal was to make sure that the beat photodiodes weren't the issue here, I think the electronics are already exonerated since driving the channel with a Marconi doesn't produce these noisy features). The EX spectrum remains noisy. I've switched everything back to the nominal configuration now to avoid further confusion. So it would appeat that this is real frequency noise that gets added in the EX fiber somehow. What can I do to fix this? The source of coupling isn't at the PSL table, else the EY channel would also show similar features. Visually, nothing seems wrong to me at EX either. So the problem is somehow in the cable tray along which the 40m of fiber is routed? This is already inside some nice foam/tubing setup, what can be done to improve it? Still doesn't explain why it suddenly became noisy...

Attachment 1: ALS_ERR_MON.pdf  48 kB  | Hide | Hide all
ALS_ERR_MON.pdf
Attachment 2: AUXnoise.pdf  188 kB  | Hide | Hide all
AUXnoise.pdf
ELOG V3.1.3-